How much 0.7MKI solution in liters will completely precipitate the Pb2+ in 2.1L of 0.18MPb(NO3)2 solution?
Answer #1

no of moles of Pb(NO3)2 = molarity * volume in L

                                         = 0.18*2.1   = 0.378moles

2KI(aq) + Pb(NO3)2(aq) -----------> PbI2(s) + 2KNO3(aq)

1 mole of Pb(NO3)2 react with 2 moles of KI

0.378 moles of Pb(NO3)2 react with = 2*0.378/1   = 0.756moles of KI

no of moles of KI   = molarity * volume in L

0.756                   = 0.7* volume in L

volume in L     = 0.756/0.7    = 1.08L

volume of KI   = 1.08L   >>>>answer